OpenStudy (anonymous):

Use Cramer's rule to solve the system. 2x + 4y - z = 32 x - 2y + 2z = -5 5x + y + z = 20 A. {( 1, -9, -6)} B. {( 2, 7, 6)} C. {( 9, 6, 9)} D. {( 1, 9, 6)}

OpenStudy (anonymous):

@Psymon

OpenStudy (nincompoop):

http://www.youtube.com/watch?v=iMQRo0tHORw

OpenStudy (uri):

lol!

OpenStudy (uri):

2x + 4y - z = 32 x - 2y + 2z = -5 5x + y + z = 20 This is the question,right?

OpenStudy (uri):